OpenStudy (anonymous):

(f(x)-f(1))/(x-1) if f(x) = ((x+3)/(x+1)), I need to reduce this and I dont know how!

OpenStudy (anonymous):

I got tired of trying to write this out sorry:)

myininaya (myininaya):

\[\frac{\frac{x+3}{x+1}-\frac{1+3}{1+1}}{x-1}=\frac{\frac{x+3}{x+1}-2}{x-1}\] \[=\frac{x+1}{x+1} \cdot \frac{\frac{x+3}{x+1}-2}{x-1}=\frac{(x+3)-2(x+1)}{(x+1)(x-1)}\] \[=\frac{-x+1}{(x-1)(x+1)}=\frac{-(x-1)}{(x-1)(x+1)}=\frac{-1}{x+1}\]

myininaya (myininaya):

and thats of course assuming \[x \neq 1 \]

OpenStudy (anonymous):

ok, Thank you so much !! this website is great, I will come here and help people myself more often

myininaya (myininaya):

and we should also assume \[x \neq -1\] since i multiplied by \[\frac{x+1}{x+1}\]

OpenStudy (anonymous):

Yeah, on the second line, but I dont quite understand why?

myininaya (myininaya):

we have a compound fraction

myininaya (myininaya):

if i multiplied the top by x+1 it would could clear the denominators on top but if i multiply by x+1 on top i must do it to the bottom

OpenStudy (anonymous):

oh, so you wanted to clean the denominator on top, I see. I though we had to flip the x-1to bring it on top, multiplying the equation by 1/ (x-1)

myininaya (myininaya):

i don't know what you are saying we could have also combined fractions on top and then change division to multiplication by fliping the second fraction

myininaya (myininaya):

\[\frac{\frac{x+3}{x+1}-2}{x-1}=\frac{\frac{x+3}{x+1}-2\frac{x+1}{x+1}}{x-1}\] \[=\frac{x+3-2(x+1)}{x+1} \div (x-1)\] \[=\frac{-x+1}{x+1} \cdot \frac{1}{x-1}=\frac{-x+1}{(x+1)(x-1)}=\frac{-(x-1)}{(x+1)(x-1)}\]

OpenStudy (anonymous):

I agree, but I though the answer was going to be 2 formulas, We initially had to determine the conditions for this function to exist. the x dosnt equal = 1 is good but I though there was something else.

OpenStudy (anonymous):

yes thats the way I though it would be. But just a second

OpenStudy (anonymous):

I guess you're right. I had x squares, in my formula, but you just canelled them out. Good so the only condition would be \[x \neq1\]

OpenStudy (anonymous):

?